# Interpreting the Weight Space of Customized Diffusion Models

### v1

@@ Abstract
Figure 1: weights2weights (w2w) space enables generative applications on model weights. We model a manifold over customized diffusion model weights as a subspace encoding different human identities. This forms a space that supports inverting the visual identity from a single image into a model, editing the identity encoded in the model, and sampling new models that encode diverse instances of people. This space effectively serves as a latent space over customized models.
@@ 1 Introduction
Creating this space unlocks a variety of applications that involve traversal in w2w (Fig. 1). First, we demonstrate that sampling model weights from w2w space corresponds to a new identity. Second, we find linear directions in this space corresponding to semantic edits in the identity. Finally, we show that enforcing weights to live in this space enables a diffusion model to learn an identity given a single image, even if it is out of distribution.
@@ 2 Related Work
Latent space of generative models. Linear latent space models of facial shape and appearance were studied extensively in the 1990s, using both PCA-based representations (e.g. Active Appearance Models [5], 3D Morphable Models [3]) as well as models operating directly in pixel and keypoint space [46]. However, these techniques were restricted to aligned and cropped frontal faces. More recently, Generative Adversarial Networks (GANs), particularly the StyleGAN series [22, 23, 24, 21], have showcased editing capabilities facilitated by their interpretable latent space. Furthermore, linear directions can be found in their latent space to conduct semantic edits by training linear classifiers or applying PCA [53, 14]. Several methods aim to project real images into the GAN latent space in order to conduct this editing [65, 1, 44, 56].
@@ 3.1 Preliminaries
In this section, we first introduce latent diffusion models (LDM) [45], which we will use to create a dataset of weights. Then, we explain the approach for deriving identity-specific models from LDM via Dreambooth [47] fine-tuning. We finally present a version of fine-tuning that uses low-dimensional weight updates (LoRA [17]). We will use the fine-tuned low-dimensional per-identity weights as data points to construct the weights manifold in Sec. 3.2.
@@
where the first term corresponds to the standard diffusion denoising objective using the subject-specific data $\mathbf{x}$ conditioned on the text prompt “[identifier] [class noun]” (e.g., “V* person”), denoted $\mathbf{c}$. The second term, weighted by $\lambda$, corresponds to a prior preservation loss, which involves the standard denoising objective using the model’s own generated samples $\mathbf{x}^{\prime}$ for the broader class $\mathbf{c}^{\prime}$ (e.g., “person”). This prevents the model from associating the class name with the specific instance, while also leveraging the semantic prior on the class. We utilize this approach to obtain a per-subject model and use its weights to create the interpretable weights manifold.
@@ 4 Experiments
We demonstrate w2w space on human identities for a variety of applications. We begin by detailing implementation details. Next, we use w2w space for 1) sampling new models encoding novel identities, 2) editing identity attributes in a consistent manner via linear traversal in w2w space, 3) embedding a new identity given a single image, and 4) projecting out-of-distribution identities into w2w space. Finally, we analyze how scaling the number of models in our dataset of model weights affects the disentanglement of attribute directions and preservation of identity.
@@ Appendix G Timestep Analysis
Edits in w2w Space correspond to identity edits with minimal interference with other visual concepts. Although not a focus, image editing is achieved as a byproduct. For further context preservation, edits in w2w Space can be integrated with delayed injection [4, 61, 11], where after $T$ timesteps, the edited weights are used instead of the original ones. We visualize this in Fig. 24. Larger $T$ in the range $[700,1000]$ are helpful for more global attribute changes, while smaller $[400,700]$ can be used for more fine-grained edits. However, by decreasing the timestep $T$, the strength of the edit is lost in favor of better context preservation. For instance, the dog’s face is better preserved in the second row at $T=600$, although the man is not as chubby compared to other $T$.

### v2

@@ Abstract
Figure 1: weights2weights (w2w) space enables generative applications on model weights. We model a manifold of customized diffusion model weights as a subspace encoding different human identities. This forms a space that supports inverting the visual identity from a single image into a model, editing the identity encoded in the model, and sampling new models that encode diverse instances of people. This space effectively serves as a latent space over customized models.
@@ 1 Introduction
Creating this space unlocks a variety of applications that involve traversal in w2w (Fig. 1). First, we demonstrate that sampling model weights from w2w space corresponds to a new identity. Second, we find linear directions in this space corresponding to semantic edits of the identity. Finally, we show that enforcing weights to live in this space enables a diffusion model to learn an identity given a single image, even if it is out of distribution.
@@ 2 Related Work
Latent space of generative models. Linear latent space models of facial shape and appearance were studied extensively in the 1990s, using both PCA-based representations (e.g. Active Appearance Models [7], 3D Morphable Models [4]) as well as models operating directly in pixel and keypoint space [49]. However, these techniques were restricted to aligned and cropped frontal faces. More recently, Generative Adversarial Networks (GANs), particularly the StyleGAN series [25, 26, 27, 24], have showcased editing capabilities facilitated by their interpretable latent space. Furthermore, linear directions can be found in their latent space to conduct semantic edits by training linear classifiers or applying PCA [56, 17], among other methods for discovering semantic directions [60, 6]. Several methods aim to project real images into the GAN latent space in order to conduct this editing [69, 1, 47, 59].
@@ 3.1 Preliminaries
In this section, we first introduce latent diffusion models (LDM) [48], which we will use to create a dataset of weights. Then, we explain the approach for obtaining identity-specific models from LDM via Dreambooth [50] fine-tuning. We finally present a version of fine-tuning that uses low-dimensional weight updates (LoRA [20]). We will use the fine-tuned low-dimensional per-identity weights as data points to construct the weights manifold in Sec. 3.2.
@@
where the first term corresponds to the standard diffusion denoising objective using the subject-specific data $\mathbf{x}$ conditioned on the text prompt “[identifier] [class noun]” (e.g., “[v] person”), denoted $\mathbf{c}$. The second term, weighted by $\lambda$, corresponds to a prior preservation loss, which involves the standard denoising objective using the model’s own generated samples $\mathbf{x}^{\prime}$ for the broader class $\mathbf{c}^{\prime}$ (e.g., “person”). This prevents the model from associating the class name with the specific instance, while also leveraging the semantic prior on the class. We utilize this approach to obtain a per-subject model and use its weights to create the interpretable weights manifold.
@@ 4 Experiments
We demonstrate w2w space on human identities for a variety of applications. We begin by detailing implementation details. Next, we use w2w space for 1) sampling new models encoding novel identities, 2) editing identity attributes in a consistent manner via linear traversal in w2w space, 3) embedding a new identity given a single image, and 4) projecting out-of-distribution identities into w2w space. Finally, we analyze how scaling the number of models in our dataset of model weights affects the disentanglement of attribute directions and preservation of identity. We refer the reader to Appendix H, where we apply w2w space to other visual concepts.
@@ Appendix G Timestep Analysis
Edits in w2w Space correspond to identity edits with minimal interference with other visual concepts. Although not a focus, image editing is achieved as a byproduct. For further context preservation, edits in w2w Space can be integrated with delayed injection [32, 65, 14, 5], where after $T$ timesteps, the edited weights are used instead of the original ones. We visualize this in Fig. 24. Larger $T$ in the range $[700,1000]$ are helpful for more global attribute changes, while smaller $[400,700]$ can be used for more fine-grained edits. However, by decreasing the timestep $T$, the strength of the edit is lost in favor of better context preservation. For instance, the dog’s face is better preserved in the second row at $T=600$, although the man is not as chubby compared to other $T$.
@@
Appendix H Extending to Other Domains
We find that similar subspaces can be created for other visual concepts beyond human identites. For instance, we apply the weights2weights framework to models encoding different dog breeds. We generate images with Stable Diffusion based on each of the $120$ dog classes from ImageNet [9]. We then conduct Dreambooth finetuning on each set of images to create a dataset of $120$ dog-encoding models. To find edit directions, we use GPT-4 [2] to create labels for each dog breed (e.g., wavy hair or not) and then train linear classifiers on the model weights like Sec. 3.3. We present examples of editing the dog-encoding models in Fig. 25. These hint at the generality of weights2weights, although more visual concepts should be explored to make a concrete conclusion.
Figure 25: Applying weights2weights edits on dog-encoding models. This suggests that weights2weights space can be constructed on domains beyond human identities.